Nightlog: 20160711 UT - D time.

Mountain: David

Tucson: Jenny, Shane, Andrew.

Skype: Olga, Barry.

17:15 LUCI1 Darks underway,

17:29 LUCI2 calibrations running in concert with LUCI1

18:04 LUCI1 darks files 0008 and 0009 show banding

19:03 LUCI cals finished, putting LUCI2 in safe state

19:06 parking LUCI2 robot

19:09 Opening. Powering down LUCI2

19:35 Sunset.

20:07 Slewing to telluric HIP 56147. Acquiring at y=580

20:16 Starting science exp. Seeing 1.1 on DIMM, 0.9 on guider.

20:20 Telluric spectrum peaks at ~17000 per read. We have plenty of time before twilight so we\x92re going to repeat with a lower exp time = 4s

20:26 Slewing to J11593.

20:32 Guide star issue, possibly too close to patrol field edge. Picking another and trying again.

20:35 Sending preset with new GS. Still no good, perhaps still too bright. Going back to original GS and waiting.

20:42 12 degree twilight.

20:54 Guide star sorted, WFSing

20:56 Fire alarm

21:09 18 degree twilight.

21:11 Observers in the lobby of the Steward building waiting for the all clear to return to the remote observing room.

21:24 Scheduler getting error to commit. Restarting.

21:32 Finally got preset sent after 2 scheduler restarts

21:33 Both guide stars we have tried are extremely dim but shouldn\x92t be. Target is getting too low, so we are moving on.

21:37 Slewing to LUCI Imaging FF1532. Star not there. David is performing an azimuth unwrap and pointing check.

21:46 Sending science preset again. Guide star is ok

21:54 Had to restart GIERS, all sorted and WFSing

22:09 About 0\x94.6 measured on image. 0.92\x94 on DIMM

22:26 Slewing to FF1725 LUCI Imaging

22:58 DIMM has not been stable for interrupt. Wind is picking up.

23:00 Reconfiguring for MODS1

23:21 Slewing to Q2048, MODS1 longslit. Pa -60

23:25 Seeing 1.0-1.1\x94 on Dimm

23:39 Offset was -4.80 +8.84

23:40 Starting science exposure

00:14 Slewing to standard HIP109452,

00:20 Had some trouble aligning, rerunning acqmods, Final pixel position 500, 619.6 with offset of -3.2\x94, 10.34\x94.

00:33 Starting telluric science

00:38 First blue image (0005) shows several bad rows near top of 2 bottom quadrants, 19 pixels wide, in 0006 the feature moved to the top 2 quadrants and is 2 pixel wide.

00:44 Slewing for science acq: Q2158 pa -46

00:58 Science started on Q22160 pa -46

1:12 Another MODSB 0008 science image with bad rows, upper quadrants.

1:25 MODS1R now showing bad rows, 0023

1:34 Slewing to V339

1:45 Single 60 second exposure to constrain scripts

1:54 Running 3x300s on V339 (peak ~40K for halpha, ~31K for OIII)

02:18 Slewing to ASN

02:28 Extreme seeing bubble. DIMM now reads > 3\x94.

02:45 Seeing still poor and variable, now ~ 2\x94. It\x92s very difficult to even clearly identify it, see MODS1 red arm file 0034. We are taking further acquisition images and monitoring the seeing.

03:04 We got a much better acquisition image! 0038. We think we can identify the science target now.

03:10 Starting science. We won\x92t be able to finish all of the exposures, but we might be lucky.

03:40 18 degree twilight.

03:44 It seems MODS1 blue exposures have stopped reading out. Running blue fitsflush.

03:48 Files recovered.

04:04 Completed 2 of 3 1200s exposures. Seeing ~ 1.3, but worse (~1.6\x94) at points. Sky bright. Closing up. We are not 100% certain the correct target was in the slit. The target wasn\x92t visible in the finding chart and there were two possible choices. We made a judgment call to proceed since the night was ending.

04:19 12 degree twilight.

04:21 LUCI1 H band imaging flats, N3.75. 10 no lamp 3.8s starting with file 122.

04:22 Starting MODS1 0\x94.8 slit flats. First b 0017, first r 0043.

04:23 LUCI1 H band imaging flats, N3.75. 10 lamp on, Halo3, 3.8s starting with file 132. ~8.4 kADU.

04:26 Taking cal unit out. Putting blind filters/mask in.

04:32 Shutting down LUCI1.

04:45 Stopping MODS1 services.

05:16 Sunrise.

-- %USERSIG{AndrewCardwell - 2016-07-11}%

Comments


 
Topic revision: r2 - 11 Jul 2016, AndrewCardwell
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback